OpenStudy (scott208):

Billy Goats invested some money in stocks and bonds. The total amount he invested was 165,000. If he invested 4.5 times as much in stocks as he did in bonds, what was his total investment in stocks?

OpenStudy (scott208):

@satellite73

OpenStudy (mathstudent55):

In problems of this type, start by letting a variable equal the simpler unknown. Let x = amount he invested in bonds. He invested 4.5 times as much in stocks than in bonds, so if x = amount invested in bonds, how would you express the amount he invested in stocks using x? Once you express that, add the two amounts and set equal to 165,000. Then solve for x.

OpenStudy (scott208):

So like. 4.5*x=165,000?

OpenStudy (scott208):

@mathmate

OpenStudy (scott208):

@Photon336

OpenStudy (mathmate):

x=amount in bonds 4.5x = amount in stocks. Total = 165000 Solve for x, and hence 4.5x

OpenStudy (scott208):

So 165,000/4.5

OpenStudy (scott208):

Which is 36,666.6 right?

OpenStudy (photon336):

I agree with @mathstudent55 \[4.5x = (stocks) \] \[x = bonds \] \[(4.5x+x) = 165,000\]

OpenStudy (scott208):

So 5.5x=165,000

OpenStudy (scott208):

And then x=30,000

OpenStudy (scott208):

So he spent 30,000 in bonds. And then we subtract that from 165,000?

OpenStudy (scott208):

So he spent 135,000 in stocks?

OpenStudy (mathmate):

Yes, that is correct. You could check using the original data: 135000+30000=165000, good 135000/30000=4.5, exactly the given ration. Final check is to make sure you have attributed the proper amount for stocks and for bonds, and I will leave that to you. You will need to reread the question to do this check.